I forgot to bring my book home and only have the problem written down.
Maragaret and Sylvia leave the house at the same time traveling in opposite directions. Margaret travels 30mph faster than sylvia. After 2 hours they are 200 miles apart. How fast is each woman traveling?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

2(30+X)+2X=200
60+2X+2X=200
4X=200-60
4X=140
X=140/4
X=35 MPH FOR SYLVIA.
35+30=65 MPH FOR MARGARET.
PROOF
2*65+2*35=200
130+70=200
200=200

Answer by stanbon(48568)

Margaret and Sylvia leave the house at the same time traveling in opposite directions. Margaret travels 30mph faster than sylvia. After 2 hours they are 200 miles apart. How fast is each woman traveling?
-------------------------------
Margaret DATA:
Distance = x miles ; Time = 2 hrs ; rate = d/t = x/2 mph
-
Sylvia DATA:
Distance = 200-x miles ; Time = 2 hrs ; Rate = d/t = (200-x)/2
----------------------------
EQUATION:
Margaret rate = Sylvia rate +30
x/2 = (200-x)/2 + 30
x = 200-x + 60
2x = 260
x = 130 miles
---------------------------
Margaret's rate = x/2 = 65 mph
===================================
Cheers,
Stan H.